OpenStudy (anonymous):

Please I really need help in Algebra 2 I WILL give medals and become a fan! what is the percent rate of change in function y= (0.98)^x determine whether the function represents exponential growth of exponential decay 2% growth 20% growth 2% decay 0.2% decay 5logb(y) + 4logb(x) logb(y^5+x^4) logb(y^5x^4 (5+4) logb(y+x) logb(yx^5+4) 3e^2x + 2 = 28 1.0797 0.4689 0.9261 2.1595 I really need help with these three problems. Thank you in advance!

OpenStudy (anonymous):

So for number 1: 2%=0.02=1-0.02=0.98, So its exponential decay 2%, which is c. Did you understand that part?

OpenStudy (anonymous):

And for 2. logby^5+logbx^4 logb(y^5x^4) so b For 3 3e^2x + 2 = 28 x=2.1595 so d.

OpenStudy (anonymous):

@zeinas. Hey thanks I really appreciate the help!

OpenStudy (anonymous):

You're welcome. :)

OpenStudy (anonymous):

@zeinas. could you help me with this one?

OpenStudy (anonymous):

5logb(y) + 4logb(x)

OpenStudy (anonymous):

Sure. It's: logby^5+logbx^4 logb(y^5x^4) It's the same as the second question before.